# Adaptive Force-Based Control of Dynamic Legged Locomotion over Uneven Terrain

### v1

@@ Abstract
Agile-legged robots have proven to be highly effective in navigating and performing tasks in complex and challenging environments, including disaster zones and industrial settings. However, these applications normally require the capability of carrying heavy loads while maintaining dynamic motion. Therefore, this paper presents a novel methodology for incorporating adaptive control into a force-based control system. Recent advancements in the control of quadruped robots show that force control can effectively realize dynamic locomotion over rough terrain. By integrating adaptive control into the force-based controller, our proposed approach can maintain the advantages of the baseline framework while adapting to significant model uncertainties and unknown terrain impact models. Experimental validation was successfully conducted on the Unitree A1 robot. With our approach, the robot can carry heavy loads (up to 50% of its weight) while performing dynamic gaits such as fast trotting and bounding across uneven terrains.
@@ I Introduction
Legged robots have numerous potential uses, from search and rescue operations to autonomous construction. To perform these tasks effectively, it is important for the robot to have an accurate understanding of the environment it will be operating in. However, due to the complexity of the robot and the environment, the model of the robot itself might contain a significant level of uncertainty and affect the robot’s stability, particularly when performing agile movements. To overcome these challenges, there is a need for the development of a control framework that can effectively compensate for these uncertainties in real-time.
@@
The utilization of convex model predictive control (MPC) with the single rigid body (SRB) model in legged robots [1] has greatly enhanced the real-time implementation of diverse walking gaits. Unlike the balance controller based on quadratic programming [2], MPC offers the capability to perform agile motions like jumping [3, 4] and high-speed bounding [5] for quadruped robots. Additionally, MPC exhibits robustness in traversing rough and uneven terrains. However, it is important to note that MPC assumes perfect knowledge of the dynamic model.
@@
To enhance trajectory tracking in the presence of unknown and changing disturbances, researchers have explored the combination of MPC with adaptive control techniques [6, 7, 8]. Additionally, parameter estimation techniques have been employed to further improve the robustness of the control system [9]. These approaches aim to adapt the controller and estimate system parameters to effectively compensate for uncertainties and disturbances, leading to improved trajectory tracking performance. It is worth noting that all of these studies were conducted using a position-based controller model.
@@
In this work, we tackle the legged robot locomotion issue in real-world scenarios with a significant level of uncertainty. The uncertainty can come from both the robot model and the environment. Since our proposed method is based on a force controller, it retains the advantage of robustness to uneven terrain. Thanks to MPC as our baseline controller, our framework can be extended to different locomotion gaits and trajectories without adjusting the controller parameters. Additionally, by incorporating the adaptive controller, our control system can handle significant model uncertainty. As a result, our approach enables legged robots to move across different terrains with unknown impact models.
@@
Figure 1: Our proposed adaptive MPC is successfully validated in an experiment on a Unitree A1 robot while carrying an unknown load of 5 kg (almost 50% of body weight) on rough terrain. Experimental results video: https://youtu.be/QmwyysdTk1k.
@@ I-A1 Offline Learning
The offline learner can either leverage a model-based control approach or learn the control system from scratch. Using a model-based method, researchers mainly target learning the dynamic to improve the controller performance [10]. One example of this approach is the integration of deep learning with MPC, in which the proposed model tries to learn the cost or dynamic terms of an MPC [11]. This hybrid method shows considerable improvement for the aerial robot [12] when learning the dynamic model from experimental data. The major limitation of this method is that it is restricted to the dynamic model learned during the training phase. However, the dynamic model is prone to frequent changes in real-world scenarios due to environmental uncertainties and external disturbances.
To overcome the limitations of previous approaches, there has been growing interest in utilizing reinforcement learning (RL) to train models from scratch. The key advantage of RL models is their ability to adapt swiftly to changes in real-world environments due to being trained in diverse environments with varying properties. In the case of quadruped robots, an RL model can directly predict appropriate joint torques for traversing different types of terrain, as demonstrated by Chen et al. [13]. Additionally, Bellegarda et al. [14] enable quadrupeds to run quickly while carrying unknown loads by training the model to learn foot positions. However, these methods heavily rely on domain randomization during training to generalize well to challenging environments. Yang et al. [15] also propose an end-to-end RL method that utilizes proprioceptive states and visual feedback to predict environmental changes.
@@ I-A2 Online Learning
To address inaccuracies in model-based controllers, researchers have explored an alternative approach using online learning, particularly supervised learning methods [16, 17, 18]. In this approach, the focus is on learning disturbances online [19], and in some cases, researchers also aim to learn the dynamics of the system itself [20]. Furthermore, this approach has been successfully applied for online calibration of kinematic parameters in legged robots [21]. In addition to that, in a recent study, a Lipschitz network method has been developed to bridge the model-reality gap in real-time [22]. The online learning method shares a close relationship with adaptive control, and numerous studies have explored the combination of these two approaches [23]. This combination aims to leverage the advantages of both methods, allowing for dynamic adaptation and continuous learning from real-time data to improve control system performance. Perhaps closest to our work in terms of online adaption is the learning method presented in [24] for legged robots. The authors correct the model behind the controller using a supervised learner while the robot is walking in an unknown environment. The data is collected during the robot’s operation to learn a linear residual model which can compensate for system errors. However, in the transition from simulation to experiment, the acceleration estimators make noisy data required for training the model. As a result, the method is only applied to estimate the linear terms since the angular terms data proved to be too noisy to be helpful in the model.
@@ I-A3 Adaptive Control
The goal of adaptive control is to tune the controller’s variables online during deployment [25]. Adaptive control has been applied for manipulation tasks to robotic arms [26], mobile robots [27, 28, 29], and quadruped robots [30, 31]. The conventional Model Reference Adaptive Control (MRAC) architecture was originally designed for controlling linear systems in the presence of parametric uncertainties [32, 33]. However, it lacks the ability to characterize the input/output performance of the system during the transient phase. To address this limitation and improve the transient performance of adaptive controllers, the $L_{1}$ adaptive control offers several advantages over traditional MRAC, such as decoupling adaptation and robustness within a control framework [34]. In addition, by incorporating a low-pass filter in adaptation law, the $L_{1}$ adaptive control can provide stability [35] and transient performance [36]. Therefore, the $L_{1}$ adaptive control technique guarantees robustness with fast adaptation [37], an essential criterion in dynamic robotics applications. Recently, by integrating $L_{1}$ adaptive controller and Bayesian learner, researchers leverage the fast adaption performance of the $L_{1}$ adaptive controllers and introduce a safe simultaneous control and learning framework [38, 39].
For legged robots, the adaptive controller has also been employed to find the value and location of the center of mass [40]. Our work on $L_{1}$ adaptive control for bipedal robots [41] considers a Control Lyapunov Function (CLF)-based controller as a closed-loop nonlinear reference model for the $L_{1}$ adaptive controller. It was validated for the robot’s walking [42] and running [43]. However, the control framework in this prior work is based on Hybrid Zero Dynamics [44], which uses joint position control to track the desired trajectory from optimization for each robot joint. Moreover, in [45], an adaptive control based on a CLF is designed for quadrupeds to interact with unknown objects. Then, they combined the criteria derived by adaptive control as a constraint in an MPC framework. However, adding more inequality constraints to MPC makes the controller more complex in terms of computation. In our approach, we compute a residual vector for compensating dynamic uncertainty, which makes the controller more time-efficient. Additionally, by employing our method, the robot is able to adapt to terrains with unknown impact models.
@@ I-B Contributions
A preliminary version of this research previously appeared in [46]; however, this paper presents several novel contributions to the prior work. This work incorporates the $L_{1}$ adaptive controller into the model predictive control (MPC). The proposed control system leverages MPC due to its robustness to uneven terrain, contact constraint, and generalization to different locomotion gaits. Moreover, by integrating adaptive control into MPC, the proposed model can compensate for significant model uncertainty. In the previous work [46], the robot can only perform quasi-static walking; however, in this work, the robot can perform dynamic motions thanks to MPC. Finally, the authors present new hardware experiments to demonstrate the effectiveness of the proposed adaptive MPC (as illustrated in Fig. 1). The main contributions of the paper are as follows:
@@
  • Thanks to MPC, our approach offers greater versatility as it can be adapted to a wide range of locomotion gaits and trajectories. Moreover, our method can handle terrain uncertainty, allowing the robot to navigate rough terrains, such as grass and gravel, as well as high-sloped terrain.
@@
  • Our proposed approach enables the control system to adapt to terrains with unknown impact models, such as soft terrain. Traversing soft terrain is a challenging task for quadruped robots. The A1 robot can walk on double-foam terrain in different directions using our method. In comparison, the robot cannot maintain its balance using the baseline controller, resulting in a collapse.
@@ II-A Gait Scheduler
In our previous work [46], we focus on the application of load-carrying tasks, where the load is unknown to the robot or the control system. Having more legs on the ground during walking could also mean that the robot could produce a larger total ground reaction force to support the heavy load. Therefore, we used a quasi-static walking gait to maximize the number of legs on the ground during walking (i.e., 3 stance legs and 1 swing leg throughout the gait). However, in this paper, our framework is not limited by any specific gait. Similar to the baseline MPC control approach [1], the approach can work for different gaits by only changing the gait definition in the gait scheduler.
@@ II-B Desired Trajectory
The desired trajectory is generated based on the robot’s velocity command. The robot operator commands $xy$-velocity and yaw rate, then $xy$-position and yaw are determined by integrating the corresponding velocity. $z$ position contains a constant value of $0.3m$, and the remaining states (roll, roll rate, pitch, pitch rate, and $z$-velocity) are always zero.
@@ II-C Single Rigid Body (SRB) Model of Robot
Due to the complexity of the legged robot, a simplified rigid-body model has been used to present the system’s dynamic. This model enables us to calculate the ground reaction forces (GRFs) in real-time. A few assumptions have been made to achieve simplified robot dynamics[1]:
Assumption 1: The robot has low inertia legs, so their effect is negligible.
@@ II-D Balance Controller
One of the baseline control approach for calculating GRFs for quadruped robots is the balance controller presented in [2] based on quadratic program (QP) solver. Based on the assumptions presented in Sec. II-C, the approximated dynamic model between the body acceleration and GRFs is as follows:
@@
where $\bm{b}_{d}$ is the desired dynamics. The idea of designing $\bm{b}_{d}$ will be elaborated in Sec. IV-A. The cost function in (II-D) includes terms that consider three goals, including (1) driving the COM position and orientation to the desired trajectories; (2) minimizing the force commands; and (3) minimizing the change of the current solution $\bm{F}^{*}$ with respect to the solution from the previous time-step, $\bm{F}^{*}_{prev}$. The priority of each goal in the cost function is defined by the weight parameters $\bm{S}\in\mathbb{R}^{6\times 6}$, $\gamma_{1}$, $\gamma_{2}$ respectively.
The constraints in the QP formulation enforce friction constraints, input saturation, and contact constraints. The constraint $\underline{\bm{d}}\leq\bm{C}\bm{F}\leq\bar{\bm{d}}$ ensures that the optimized forces lie inside the friction pyramid and the normal forces stay within a feasible range. More details can be found in [2]. Besides the friction constraint, we will enforce the force constraints for the swing legs, $\bm{F}_{swing}=\bm{0}$. The swing legs are then kept in the posing position until it switches to the stance phase. More details on swing leg control are provided in Sec. II-F.
@@ II-F Swing Leg Control
For the swing legs, the final footstep location for each leg is calculated from the corresponding hip location using a linear combination of Raibert heuristic [51], and a feedback term from the capture point formulation [52, 47]. The final footstep locations ($\bm{p}_{f,i}$) are projected on an assumed ground plane and are calculated by:
@@ III Overview of the Proposed Approach
This section will present an overview of our novel control architecture to incorporate adaptive control into the force control framework. While our approach is not limited to any specific adaptive control approach, we decide to use $L_{1}$ adaptive control [37, 41] thanks to its advancement in guaranteeing fast adaptation and smooth control signals. Note that our proposed control system is designed for the stance leg control part in the control architecture of the quadruped robot (see Fig. 2).
@@
In the following sections, we will elaborate on integrating two different force-based control as the baseline controller into the adaptive control. First, in Sec. IV, we will describe the proposed method using a QP-based balancing controller, as presented in Fig. 3b. Then, in Sec. V, we will show how to incorporate MPC into the adaptive controller in detail, as illustrated in Fig. 3c.

### v2

@@ Abstract
Agile-legged robots have proven to be highly effective in navigating and performing tasks in complex and challenging environments, including disaster zones and industrial settings. However, these applications commonly require the capability of carrying heavy loads while maintaining dynamic motion. Therefore, this paper presents a novel methodology for incorporating adaptive control into a force-based control system. Recent advancements in the control of quadruped robots show that force control can effectively realize dynamic locomotion over rough terrain. By integrating adaptive control into the force-based controller, our proposed approach can maintain the advantages of the baseline framework while adapting to significant model uncertainties and unknown terrain impact models. Experimental validation was successfully conducted on the Unitree A1 robot. With our approach, the robot can carry heavy loads (up to 50% of its weight) while performing dynamic gaits such as fast trotting and bounding across uneven terrains.
@@ I Introduction
Legged robots have numerous potential uses, from search and rescue operations to autonomous construction. To perform these tasks effectively, the robot must accurately understand the environment it will be operating in. However, due to the complexity of the robot and the environment, the robot’s model might contain a significant level of uncertainty and affect the robot’s stability, particularly when performing agile movements. To overcome these challenges, there is a need to develop a control framework that can effectively compensate for these uncertainties in real-time.
@@
The utilization of convex model predictive control (MPC) with the single rigid body (SRB) model in legged robots [1] has greatly enhanced the real-time implementation of diverse walking gaits. Unlike the balance controller based on quadratic programming [2], MPC offers the capability to perform agile motions like jumping [3, 4] and high-speed bounding [5] for quadruped robots. Additionally, MPC exhibits robustness when traversing rough and uneven terrains. However, it is important to note that MPC assumes perfect knowledge of the dynamic model.
@@
To enhance trajectory tracking in the presence of unknown and changing disturbances, researchers have explored the combination of MPC with adaptive control techniques [6, 7, 8]. Additionally, parameter estimation techniques have been employed to improve the robustness of the control system further [9]. These approaches aim to adapt the controller and estimate system parameters to effectively compensate for uncertainties and disturbances, leading to improved trajectory tracking performance. It is worth noting that all of these studies were conducted using a position-based controller model.
@@
In this work, we tackle the legged robot locomotion issue in real-world scenarios with significant uncertainty. The uncertainty can come from both the robot model and the environment. Since our proposed method is based on a force controller, it retains the advantage of robustness to uneven terrain. Thanks to MPC as our baseline controller, our framework can be extended to different locomotion gaits and trajectories without adjusting the controller parameters. Moreover, in our control system, we effectively manage substantial model uncertainty by utilizing the adaptive controller. By implementing adaptive control, our framework evolves into a versatile solution for mitigating persistent disturbances across various operations and over time. Given the adaptive control’s capability to address uncertainties continuously, it provides a practical approach for real-world applications in legged robot autonomy, such as rescue missions, inspections, and logistics. This ability to compensate for persistent disturbances in real-world scenarios eliminates the need for recalibration for various tasks, enabling a thorough online operation. As a result, this represents a key contribution to our work, offering a comprehensive approach for legged robot applications and facilitating movement across diverse terrains with unknown impact models.
@@
Figure 1: Our proposed adaptive MPC is successfully validated in an experiment on a Unitree A1 robot while carrying an unknown load of 5 kg (almost 50% of body weight) on rough terrain. Experimental results video: https://youtu.be/5t1mSh0q3lk.
@@ I-A1 Offline Learning
The offline learner can leverage a model-based control approach or learn the control system from scratch. Using a model-based method, researchers mainly target learning the dynamic to improve the controller performance [10]. One example of this approach is integrating deep learning with MPC, in which the proposed model tries to learn the cost or dynamic terms of an MPC [11]. This hybrid method shows considerable improvement for the aerial robot [12] when learning the dynamic model from experimental data. The major limitation of this method is that it is restricted to the dynamic model learned during the training phase. However, the dynamic model is prone to frequent changes in real-world scenarios due to environmental uncertainties and external disturbances.
There has been growing interest in utilizing reinforcement learning (RL) to train models from scratch to overcome the limitations of previous approaches. The key advantage of RL models is their ability to adapt swiftly to changes in real-world environments due to being trained in diverse environments with varying properties. In the case of quadruped robots, an RL model can directly predict appropriate joint torques for traversing different types of terrain, as demonstrated by Chen et al. [13]. Additionally, by training the model to learn foot positions, Bellegarda et al. [14] enable quadrupeds to run quickly while carrying unknown loads. However, these methods heavily rely on domain randomization during training to generalize to challenging environments. Yang et al. [15] also propose an end-to-end RL method that utilizes proprioceptive states and visual feedback to predict environmental changes.
@@ I-A2 Online Learning
To address inaccuracies in model-based controllers, researchers have explored an alternative approach using online learning, mainly supervised learning methods [16, 17, 18]. In this approach, the focus is on learning disturbances online [19], and in some cases, researchers also aim to learn the dynamics of the system itself [20]. Furthermore, this approach has been successfully applied for online calibration of kinematic parameters in legged robots [21]. In addition, a recent study has developed a Lipschitz network method to bridge the model-reality gap in real-time [22]. The online learning method is closely related to adaptive control, and numerous studies have explored combining these two approaches [23]. This combination aims to leverage the advantages of both methods, allowing for dynamic adaptation and continuous learning from real-time data to improve control system performance. Perhaps closest to our work in terms of online adaption is the learning method presented in [24] for legged robots. The authors correct the model behind the controller using a supervised learner while the robot is walking in an unknown environment. The data is collected during the robot’s operation to learn a linear residual model that can compensate for system errors. However, in the transition from simulation to experiment, the acceleration estimators make noisy data required for training the model. As a result, the method is only applied to estimate the linear terms since the angular terms data proved to be too noisy to be helpful in the model.
To enhance controller efficiency and performance, auto-tuning methods, particularly for PID controllers, have gained widespread use [25, 26]. These methods fall into two categories: model-based and model-free. Model-based approaches use system model information, often employing the gradient of the performance criterion to enhance local performance [27]. In contrast, model-free methods, such as Markov chain Monte Carlo [28], Gaussian process [29, 30], and deep neural network [31], approximate gradients or surrogate models to boost performance. However, model-based approaches may fail in real-world scenarios due to imperfect dynamic knowledge, and model-free methods like Bayesian optimization can be inefficient in high-dimensional parameter tuning. Addressing this, recent works [32, 33] directly obtain the gradient of the loss function with respect to controller parameters and apply it to gradient descent for performance improvement. Auto-tuning generally requires data or a fixed model for training and, therefore, does not fit well for real-time and fast adaptation to significant model uncertainty.
@@ I-A3 Adaptive Control
Adaptive control aims to tune the controller’s variables online during deployment [34]. Adaptive control has been applied for manipulation tasks to robotic arms [35], mobile robots [36, 37, 38], and quadruped robots [39, 40]. The conventional Model Reference Adaptive Control (MRAC) architecture was initially designed for controlling linear systems in the presence of parametric uncertainties [41, 42]. However, it cannot characterize the input/output performance of the system during the transient phase. To address this limitation and improve the transient performance of adaptive controllers, the $L_{1}$ adaptive control offers several advantages over traditional MRAC, such as decoupling adaptation and robustness within a control framework [43]. In addition, incorporating a low-pass filter in adaptation law allows the $L_{1}$ adaptive control to provide stability [44] and transient performance [45]. Therefore, the $L_{1}$ adaptive control technique guarantees robustness with fast adaptation [46], an essential criterion in dynamic robotics applications. Recently, by integrating $L_{1}$ adaptive controller and Bayesian learner, researchers leverage the fast adaption performance of the $L_{1}$ adaptive controllers and introduce a safe simultaneous control and learning framework [47, 48].
For legged robots, the adaptive controller has also been employed to find the value and location of the center of mass [49]. Our work on $L_{1}$ adaptive control for bipedal robots [50] considers a Control Lyapunov Function (CLF)-based controller as a closed-loop nonlinear reference model for the $L_{1}$ adaptive controller. It was validated for the robot’s walking [51] and running [52]. However, the control framework in this prior work is based on Hybrid Zero Dynamics [53], which uses joint position control to track the desired trajectory from optimization for each robot joint. Moreover, in [54], an adaptive control based on a CLF is designed for quadrupeds to interact with unknown objects. Then, they combined the criteria derived by adaptive control as a constraint in an MPC framework. However, adding more inequality constraints to MPC makes the controller more complex in terms of computation. In our approach, we compute a residual vector to compensate for dynamic uncertainty, which makes the controller more time-efficient. Additionally, by employing our method, the robot can adapt to terrains with unknown impact models.
@@ I-B Contributions
A preliminary version of this research appeared in [55]; however, this paper presents several novel contributions to the prior work. This work incorporates the $L_{1}$ adaptive controller into the model predictive control (MPC). The proposed control system leverages MPC due to its robustness to uneven terrain, contact constraints, and generalization to different locomotion gaits. Moreover, by integrating adaptive control into MPC, the proposed model can compensate for significant model uncertainty. In the previous work [55], the robot can only perform quasi-static walking; however, in this work, the robot can perform dynamic motions thanks to MPC. Finally, the authors present new hardware experiments to demonstrate the effectiveness of the proposed adaptive MPC (as illustrated in Fig. 1). The main contributions of the paper are as follows:
@@
  • Thanks to MPC, our approach offers greater versatility as it can be adapted to a wide range of locomotion gaits and trajectories. Moreover, our method can handle terrain uncertainty, allowing the robot to navigate rough terrains and high-sloped terrain, such as grass and gravel.
@@
  • Our proposed approach enables the control system to adapt to terrains with unknown impact models, such as soft terrain. Traversing soft terrain is a challenging task for quadruped robots. Using our method, the A1 robot can walk on double-foam terrain in different directions. In comparison, the robot cannot maintain its balance using the baseline controller, resulting in a collapse.
@@ II-A Gait Scheduler
In our previous work [55], we focus on the application of load-carrying tasks, where the load is unknown to the robot or the control system. Having more legs on the ground during walking could also mean that the robot could produce a more significant total ground reaction force to support the heavy load. Therefore, we used a quasi-static walking gait to maximize the number of legs on the ground during walking (i.e., three stance legs and one swing leg throughout the gait). However, in this paper, our framework is not limited by any specific gait. Similar to the baseline MPC control approach [1], the approach can work for different gaits by only changing the gait definition in the gait scheduler.
@@ II-B Desired Trajectory
The desired trajectory is generated based on the robot’s velocity command. The robot operator commands $xy$-velocity and yaw rate, and then $xy$-position and yaw are determined by integrating the corresponding velocity. $z$ position contains a constant value of $0.3m$, and the remaining states (roll, roll rate, pitch, pitch rate, and $z$-velocity) are always zero.
@@ II-C Single Rigid Body (SRB) Model of Robot
Due to the complexity of the legged robot, a simplified rigid-body model has been used to present the system’s dynamic. This model lets us calculate the ground reaction forces (GRFs) in real-time. A few assumptions have been made to achieve simplified robot dynamics[1]:
Assumption 1: The robot has low inertia legs, so their effect is negligible on the robot’s rigid body dynamic.
@@ II-D Balance Controller
One of the baseline control approaches for calculating GRFs for quadruped robots is the balance controller presented in [2] based on a quadratic program (QP) solver. Based on the assumptions presented in Sec. II-C, the approximated dynamic model between the body acceleration and GRFs is as follows:
@@
where $\bm{b}_{d}$ is the desired dynamics. The idea of designing $\bm{b}_{d}$ will be elaborated in Sec. IV-A. The cost function in (II-D) includes terms that consider three goals, including (1) driving the COM position and orientation to the desired trajectories, (2) minimizing the force commands, and (3) minimizing the change of the current solution $\bm{F}^{*}$ with respect to the solution from the previous time-step, $\bm{F}^{*}_{prev}$. The priority of each goal in the cost function is defined by the weight parameters $\bm{S}\in\mathbb{R}^{6\times 6}$, $\gamma_{1}$, $\gamma_{2}$ respectively.
The constraints in the QP formulation enforce friction constraints, input saturation, and contact constraints. The constraint $\underline{\bm{d}}\leq\bm{C}\bm{F}\leq\bar{\bm{d}}$ ensures that the optimized forces lie inside the friction pyramid and the normal forces stay within a feasible range. More details can be found in [2]. Besides the friction constraint, we will enforce the force constraints for the swing legs, $\bm{F}_{swing}=\bm{0}$. The swing legs are then kept in the posing position until they switch to the stance phase. More details on swing leg control are provided in Sec. II-F.
@@ II-F Swing Leg Control
For the swing legs, the final footstep location for each leg is calculated from the corresponding hip location using a linear combination of Raibert heuristic [60] and a feedback term from the capture point formulation [61, 56]. The final footstep locations ($\bm{p}_{f,i}$) are projected on an assumed ground plane and are calculated by:
@@ III Overview of the Proposed Approach
This section will present an overview of our novel control architecture to incorporate adaptive control into the force control framework. While our approach is not limited to any specific adaptive control approach, we decided to use $L_{1}$ adaptive control [46, 50] thanks to its advancement in guaranteeing fast adaptation and smooth control signals. Note that our proposed control system is designed for the stance leg control part in the control architecture of the quadruped robot (see Fig. 2).
@@
In the following sections, we will elaborate on integrating two different force-based controls as the baseline controller into the adaptive control. First, in Sec. IV, we will describe the proposed method using a QP-based balancing controller, as presented in Fig. 3b. Then, in Sec. V, we will show how to incorporate MPC into the adaptive controller in detail, as illustrated in Fig. 3c.